# Model-Agnostic Materials Data Extraction from Text Using Language Models

### v2

@@ I Introduction
Obtaining reliable and comprehensive materials data is crucial for many research and industrial applications. If necessary information is not accessible through curated databases researchers typically must manually extract the data from research papers, a process that can be time-consuming and labor-intensive. Natural language processing (NLP) with general Language Models (LMs), and in particular, powerful large LMs (LLMs) trained on massive bodies of data, offer a new and potentially transformative technology for increasing the efficiency of extracting data from papers. These methods are particularly valuable when the data is embedded in the text of the documents, rather than being presented in a structured or tabular format, making it harder to find and extract. The rapid pace of development in NLP and frequent release of improved LLMs suggests they can be best utilized by methods which are easily adapted to new LLMs. In this paper we present such a flexible method for materials data extraction and demonstrate that it can achieve excellent precision and recall.
@@
So far, the majority of materials data extraction approaches focus on fully automatic data extraction [1, 2, 3, 4, 5, 6]. Automation is clearly desirable, particularly when extracting very large databases. However, more automation tends to require more complexity in the software, sophistication in training schemes, and knowledge about the extracted property. In addition, if a high level of completeness is required from a database, the recall of these approaches may not be sufficient. In such fully automated approaches a large amount of focus has been placed on the complex task of named entity recognition (NER) [7, 8, 9, 10, 11], so that the property, material, values and units can be extracted accurately. However, automatic identification of an improper recognition is still very challenging, which can reduce the precision of such approaches. Tools for automatic materials and chemistry data extraction, like OSCAR4 [12] or ChemDataExtractor [13, 14], have been developed and used to successfully extract large databases. A recent example includes a database of over 22 thousand entries for relatively complex thermoelectric properties [15], at an average precision of 82.5% and a recall of 39.23%, or over 100 thousand band gap values [16], with an average precision of 84% and a recall of 64%. More complex information such as synthesis recipes [17, 18, 19, 20, 21, 22] have also been extracted with automated NLP-based methods. Although not complete due to the relatively low recall, databases of that size are useful for training machine learning models [23, 24, 25, 26, 27, 28, 29], and would be very time consuming or impossible to extract with virtually any other method than full automation. Other recent examples of databases created in a similar way include photovoltaic properties and device material data for dye-sensitized solar cells [30], yield strength and grain size [31], and refractive index [32, 33]. Other notable databases gathered with NLP-based approaches include more complex information than just data values, such as synthesis procedures [18, 26]. Recently, another method for structured information extraction, making use of the GPT-3 capabilities was presented [34]. In that work, the focus is placed on the complicated NER tasks and relation extraction, at which GPT-3 excels. In that approach, more complex sentences can be successfully parsed into structured information. A ”human in the loop” approach was use to fine-tune the model, a technique that seems to be emerging as a method of choice to obtain higher performing models. Impressive performance was achieved in this work for structured information extraction, although at a price of a relatively large set of relatively complex training examples.
@@
Here $A$ is up-front fixed time to develop the method for a specific case and $B$ is a marginal time cost whose contribution scales with some measure of data quantity NN (NN represents some function of the number of papers, sentences, and data to extract). The other axis is database quality, generally represented by some combination of precision (what fraction of the found data are right) and recall (what fraction of the available data was found) of the database. A schematic plot representing required human time as a function of the size of the database for different methods, and the quality of their results, is presented in Fig. 1, where a logarithmic scale was used to emphasize the behavior for small-moderate sized databases, where the method is the most advantageous. While the classification of a dataset’s size depends significantly on the context and scientific field, in this paper we base this classification, as detailed in the following paragraph, on an evaluation of the typical quantities of relevant materials data found in existing literature. One limiting case, which we will call ”small data”, is one where only a small amount of data, up to around 100 points, is available in the literature (for example, properties that are very new, very hard to measure, or studied by only a small community), and where completeness and accuracy are highly valued. For example, as of this writing ”small data” might refer to superconductors with $T_{c}>200K$ [35]. It is typical to gather data for properties in the very small data limit fully manually, usually by experts in the field. Full manual curation is practical due to the limited number of papers and data and assures that the data is comprehensive and accurate. This fully manual approach is represented in Fig. 1 as the blue line, which is preferred (dashed) for very small amounts of available data. Even though it is technically slower than other methods, even in the very small data range, it is still the method of choice due to the highest possible quality of the results.
@@
The opposite limiting case, which we will call ”large data”, is when there is a lot of data in the literature, more than a couple thousand datapoints, the database is expected to be large, and modest precision and recall are acceptable. For example, such a database might be pursued for use for building machine learning regression and classification models on widely studied properties. For this large data case fully automated NLP-based approaches may be the most appropriate solution (see green curve in Fig. 1 for large number of entries). However, such an automated approach can result in an incomplete database that may not be sufficient for certain research or industrial applications, e.g., where extremes of performance of just a few materials might be the primary interest. In addition, fully automated approaches often require extensive retraining and building of parsers specific for different properties, as well as a significant amount of coding. These methods thus often require a significant initial investment of human time.
@@
The logic of the best approaches for these extremes is simple. Large data (e.g., $10^{4}$ data points) can be most efficiently extracted by spending human time on automating the extraction (leading to large $A$ and small $B$ in Eq. 1), and reduced precision and recall is often of limited consequence since so much data is available. Small data sets (e.g, <10<~10 data points) can be most efficiently extracted by spending human time on directly extracting the data (leading to small $A$ and large $B$ in equation Eq. 1), and high precision and recall is typically more important for smaller databases. However, the optimal approach for the middle ground between these scales, which represents many databases in materials, is not obvious.
@@
In this paper we demonstrate the method by developing databases with multiple LLMs. The simplicity and flexibility of the method is illustrated by repeating the development of a benchmark bulk modulus sentence classification database with multiple OpenAI GPT models, including the recently released GPT-3.5 davinci , GPT-3.5 (turbo) and GPT-4 [38, 39], as well as the bart- and DeBERTaV3-based language models [40, 41, 42, 43] hosted on huggingface, currently the most downloaded models for text classification. It is important to demonstrate the applicability and efficiency of the method on both simple, free, and accessible LLMs that can be easily used on a personal computer, and on LLMs which require significantly more computation and may be out of reach of some people’s resources for now. Even though there exist fully free and open LLMs, such as OPT[44], BLOOM[45] or LLaMA[46], their use is computationally expensive and not convenient, which contradicts the spirit of ease and accessibility of the presented method. Therefore, we opted for The OpenAI’s models whose API allows one to efficiently use the LLM on outside servers, although is not free. GPT-3/3.5/4 are also currently the most popular LLMs, so its a choice that will likely be relevant for many users.
@@ II Description of the method
It is worth noting that some models, such as GPT-3/3.5 davinci and GPT-3.5/4 (chat) at the time of writing this article are not free to use. . Therefore, the flexibility to use different LLMs within the method is very valuable, as some free models, while not necessarily capable of accurately performing the more complex tasks such as automated data structurization, and although overall generally less capable than GPT-based models, perform well enough in the simple task of classification to produce satisfying results. However, in the case of OpenAI GPT-3, both model usage and fine-tuning is done on outside servers, so in a situation where computational resources are not available to run locally, it may enable one to use the best models at a low cost.
@@
This optional step is just a chance for the human user to provide confirmation or correction to particularly important zero-shot classification data from Step 1 and then use those checks to fine-tune the LLM. Similar steps are often taken in other data extraction approaches, and machine learning in general [34, 36, 53]. Specifically, as the highest scoring sentences are being manually verified, a new training sets consisting of true positive and true negative examples is built. Since the precision of results of Step 1 is typically around 50% at 90% recall (see Fig. 3 (d)), the created sets are typically close to equal in size. The human labeled sets consist of positive cases, which represent true positives from Step 1, and negative cases, consisting of false positives from Step 1. The latter are the most valuable counter-examples for the negative training set, as these are the sentences the easiest for the model to confuse for positives. If after reaching the desired amount of verified positive sentences the corresponding set of negative sentences is smaller, it may be complemented with random sentences from the analyzed papers (the exceeding majority of which are negative). Fig. 4 shows how the classification model improves when fine-tuned on datasets of increasing size. A detailed analysis of that figure is present in Sec. III, where a conclusion is made that after around 100 positive sentences for the quicker learning models such as GPT-3/3.5 davinci or bart, we start to observe diminishing returns with this human labeled dataset size increase and it may not be worth spending more human time on obtaining more examples. Therefore, we recommend to perform the manual verification of the zero-shot classification until 100 positive sentences (and a corresponding 100 negative - made easy due to the  50% precision) are obtained, a number easy to remember and satisfactory for an efficient fine-tuning dataset.
This steps usually takes no more than 30 minutes for approximately 100 sentences - each sentence has to be classified only in a binary fashion which is a very simple task and takes just a few seconds per sentence. The classification is as simple and straightforward as assigning $1$ for positive and 0 to each sentence in a spreadsheet. The fine-tuning itself, for the small locally hosted models (bart and DeBERTaV3), takes around 30 minutes on a single modern CPU (Intel(R) Xeon(R) CPU E5-2670, 166 GFLOPS) and can be reduced to just a few minutes if GPUs are used. The OpenAI models are fine-tuned on external OpenAI servers in less than 30 minutes and do not require any local resources. After this step is performed and the sentences are once again reclassified using the now fine-tuned model the precision and recall are greatly improved, as can be seen in Fig. 4.
@@
Data structurization. In this step, extraction of the structured data is performed. Here, by structured data we mean the full information necessary to provide a datapoint: the name of the material/system, the numerical value of the given property, its unit, and in some cases an additional detail, such as the temperature at which the value was obtained. At the same time as the data is extracted the precision of the result is improved to perfect (or near perfect, depending on the expertise and accuracy of the human supervising this step), by simply ignoring the false positive sentences left over from previous steps. The result of this step is a final, curated structured database.
@@
The user will typically perform this step by first ranking the sentences by their probability of being relevant, which is the output from Step 1 (or Step 2 if performed), and start reviewing the list at the top, working down until they decide to stop. As the user works through the results in that fashion, they traverse down the precision recall curve (PRC) (see. Figs. 3 (a) and 4 (a)). While the recall is impossible to assess without knowing the ground truth, the user is fully aware of the precision of the data they have already analyzed, therefore using the PRC they can estimate the recall and decide to stop when a desired recall is reached (with the assumption that the PRC are similar to those shown in Figs. 3 (a) and 4 (a)). For best models, reaching recall of around 90% (close to that of a fully manual data curation) without performing the optional Step 2 happens for a precision close to 60%, while for a fine-tuned model (with Step 2), for a precision over 80%. The 90% recall seems to be a reasonable value to stop the process, as the precision sharply drops for higher values, which diminishes returns for the human time involved. However, this behavior may vary depending on the case, which will be discussed further in the Sec. III.
@@
In general, human assisted data structurization, even when only the sentences containing the relevant data are given, may be a tedious and time consuming task. However, at this point it is the only method that can guarantee an almost perfect precision. For an inexperienced user, extracting one datapoint from a given sentence and its surrounding context fully by hand may take as long as 30 seconds, depending on the complexity of the property being analyzed and how it is typically expressed in research papers. Considering this, only relatively modest sized databases are reasonable to create. However, with experience, this time quickly reduces as the user gets used to the process. In addition, more experienced users may employ simple computer codes, e.g. based on regular expressions, which would preselect possible candidates for values and units, reducing the time significantly. In the longer term, it is likely the NLP tools will help automate this data structurization step, although they do not do this very effectively at present without either human supervision or a major effort to tune them. For example, GPT-3/3.5/4 is capable of parsing unstructured data in a zero-shot fashion, with no need for retraining. In the case of our bulk modulus sentences dataset we found that in over 60% of cases GPT-3/3.5/4 is capable of correctly providing the entire data entry for a given property (name of the material/system, value, unit), and an incomplete datapoint (wrong material/system, but correct value and unit) in over 95% of cases. The only drawback that prohibits a full automation of this step with a LLM is the inability to automatically and unambiguously distinguish between correct and incorrect extracted datapoints. Even though the model does not tend to make up (hallucinate) data, it sometimes provides an incomplete or inaccurate extraction (e.g. ”alloy” instead of ”AlCu alloy” for the material, or ”100” instead of ”greater than />/> 100” for the value, etc.). However, human assistance in determining whether the data has been structurized properly, and in case it was not, fixing it by hand, can easily remedy that problem. Since almost all values and units are extracted correctly, and only less than half of the material names require fixing, using a LLM approach greatly reduces the human time and effort required for data structurization. Using an LLM we found that the average human time required to extract each good datapoint was reduced to under 10 seconds, keeping the same, almost perfect precision. Thus, and NLP-assisted data structurization, while still a tedious process, enables one create databases of up to around 1000 entries (more or less, depending on the users predisposition to and efficiency at repetitive tasks), in one workday. This timing includes the whole process, beginning (Step 0) to end (structurized database after Step 3), although almost all the human time is spent in Step 3.

### v3

# Flexible, Model-Agnostic Method for Materials Data Extraction from Text Using General Purpose Language Models

@@ I Introduction
Obtaining reliable and comprehensive materials data is crucial for many research and industrial applications. If necessary information is not accessible through curated databases researchers typically must manually extract the data from research papers, a process that can be time-consuming and labor-intensive. Natural language processing (NLP) with general Language Models (LMs), and in particular, powerful large LMs (LLMs) trained on massive bodies of data, offer a new and potentially transformative technology for increasing the efficiency of extracting data from papers. These methods are particularly valuable when the data is embedded in the text of the documents, rather than being presented in a structured or tabular format, making it harder to find and extract. While currently LLMs often fall short in practical applications, struggling with comprehending and reasoning over complex, interconnected knowledge domains, they offer significant potential for innovation in materials science and are likely to play a crucial role in materials data extraction [1].
The rapid pace of development in NLP and frequent release of improved LLMs suggests they can be best utilized by methods which are easily adapted to new LLMs. In this paper we present such a flexible method for materials data extraction and demonstrate that it can achieve excellent precision and recall.
@@
So far, the majority of materials data extraction approaches focus on fully automatic data extraction [2, 3, 4, 5, 6, 7]. Automation is clearly desirable, particularly when extracting very large databases. However, more automation tends to require more complexity in the software, sophistication in training schemes, and knowledge about the extracted property. In addition, if a high level of completeness is required from a database, the recall of these approaches may not be sufficient. In such fully automated approaches a large amount of focus has been placed on the complex task of named entity recognition (NER) [8, 9, 10, 11, 12], so that the property, material, values and units can be extracted accurately. However, automatic identification of an improper recognition is still very challenging, which can reduce the precision of such approaches. Tools for automatic materials and chemistry data extraction, like OSCAR4 [13] or ChemDataExtractor [14, 15], have been developed and used to successfully extract large databases. A recent example includes a database of over 22 thousand entries for relatively complex thermoelectric properties [16], at an average precision of 82.5% and a recall of 39.23%, or over 100 thousand band gap values [17], with an average precision of 84% and a recall of 64%. More complex information such as synthesis recipes [18, 19, 20, 21, 22, 23] have also been extracted with automated NLP-based methods. Although not complete due to the relatively low recall, databases of that size are useful for training machine learning models [24, 25, 26, 27, 28, 29, 30], and would be very time consuming or impossible to extract with virtually any other method than full automation. Other recent examples of databases created in a similar way include photovoltaic properties and device material data for dye-sensitized solar cells [31], yield strength and grain size [32], and refractive index [33, 34]. Other notable databases gathered with NLP-based approaches include more complex information than just data values, such as synthesis procedures [19, 27]. Recently, another method for structured information extraction, making use of the GPT-3 capabilities was presented [35]. In that work, the focus is placed on the complicated NER tasks and relation extraction, at which GPT-3 excels. In that approach, more complex sentences can be successfully parsed into structured information. A ”human in the loop” approach was used to fine-tune the model, a technique that seems to be emerging as a method of choice to obtain higher performing models. Impressive performance was achieved in this work for structured information extraction, although at a price of a relatively large set of relatively complex training examples.
In addition, the emergence of highly specialized LLMs underscores the rapid advancement in the field. In Ref. [36] an instruction-based process specifically designed for materials science enhanced the accuracy and relevance of data extraction. Such specialized fine-tuning shows significant advantages in dealing with niche materials science tasks.
Recently, fully automated agent-based LLM approaches to analyze scientific text have been proposed as well, which are capable of answering science questions with information from research papers [37], and generating customizable datasets [38]. Other fully automated LLM-based methods, including those that leverage complex prompt engineering workflows within LLMs have been proposed to curate large materials datasets of a higher quality than conventional automated NLP methods when used with state-of-the-art LLMs [39].
@@
Here $A$ is up-front fixed time to develop the method for a specific case and $B$ is a marginal time cost whose contribution scales with some measure of data quantity $N$ ($N$ represents some function of the number of papers, sentences, and data to extract). The other axis is database quality, generally represented by some combination of precision (what fraction of the found data are right) and recall (what fraction of the available data was found) of the database. A schematic plot representing required human time as a function of the size of the database for different methods, and the quality of their results, is presented in Fig. 1, where a logarithmic scale was used to emphasize the behavior for small-moderate sized databases, where the method is the most advantageous. While the classification of a dataset’s size depends significantly on the context and scientific field, in this paper we base this classification, as detailed in the following paragraph, on an evaluation of the typical quantities of relevant materials data found in existing literature.
One limiting case, which we will call ”small data”, is one where only a small amount of data, up to around 100 points, is available in the literature (for example, properties that are very new, very hard to measure, or studied by only a small community), and where completeness and accuracy are highly valued. For example, as of this writing ”small data” might refer to superconductors with $T_{c}>200K$ [40]. It is typical to gather data for properties in the very small data limit fully manually, usually by experts in the field. Full manual curation is practical due to the limited number of papers and data and assures that the data is comprehensive and accurate. This fully manual approach is represented in Fig. 1 as the blue line, which is preferred (dashed) for very small amounts of available data. Even though it is technically slower than other methods, even in the very small data range, it is still the method of choice due to the highest possible quality of the results.
@@
The opposite limiting case, which we will call ”large data”, is when there is a lot of data in the literature, more than a couple thousand datapoints, the database is expected to be large, and modest precision and recall are acceptable. For example, such a database might be pursued for use for building machine learning regression and classification models on widely studied properties. For this large data case fully automated NLP-based approaches may be the most appropriate solution (see green curve in Fig. 1 for large number of entries). However, such an automated approach can result in an incomplete database that may not be sufficient for certain research or industrial applications, e.g., where extremes of performance of just a few materials might be the primary interest. In addition, conventional, fully automated NLP approaches often require extensive retraining and building of parsers specific for different properties, as well as a significant amount of coding. These methods thus often require a significant initial investment of human time.
@@
The logic of the best approaches for these extremes is simple. Large data (e.g., $>~{}10^{4}$ data points) can be most efficiently extracted by spending human time on automating the extraction (leading to large $A$ and small $B$ in Eq. 1), and reduced precision and recall is often of limited consequence since so much data is available. Small data sets (e.g, $<~{}10$ data points) can be most efficiently extracted by spending human time on directly extracting the data (leading to small $A$ and large $B$ in equation Eq. 1), and high precision and recall is typically more important for smaller databases. However, the optimal approach for the middle ground between these scales, which represents many databases in materials, is not obvious.
@@
In this paper we demonstrate the method by developing databases with multiple LLMs. The simplicity and flexibility of the method is illustrated by repeating the development of a benchmark bulk modulus sentence classification database with multiple OpenAI GPT models, including the recently released GPT-3.5 davinci, GPT-3.5 (turbo) and GPT-4 [43, 44], as well as the bart- and DeBERTaV3-based language models [45, 46, 47, 48] hosted on huggingface, currently the most downloaded models for text classification. It is important to demonstrate the applicability and efficiency of the method on both simple, free, and accessible LLMs that can be easily used on a personal computer, and on LLMs which require significantly more computation and may be out of reach of some most people’s resources for now. Even though there exist fully free and open LLMs, such as OPT[49], BLOOM[50] or LLaMA[51], their use is computationally expensive and not convenient, which contradicts the spirit of ease and accessibility of the presented method. Therefore, we opted for The OpenAI’s models whose API allows one to efficiently use the LLM on outside servers, although is not free. GPT-3/3.5/4 are also currently the most popular LLMs, so its a choice that will likely be relevant for many users.
@@ II Description of the approach
It is worth noting that some models, such as GPT-3/3.5 davinci and GPT-3.5/4 (chat) at the time of writing this article are not free to use. Therefore, the flexibility to use different LLMs within the method is very valuable, as some free models, while not necessarily capable of accurately performing the more complex tasks such as automated data structurization, and although overall generally less capable than GPT-based models, perform well enough in the simple task of classification to produce satisfying results. However, in the case of OpenAI GPT-3, both model usage and fine-tuning is done on outside servers, so in a situation where computational resources are not available to run locally, it may enable one to use the best models at a low cost.
@@
This optional step is just a chance for the human user to provide confirmation or correction to particularly important zero-shot classification data from Step 1 and then use those checks to fine-tune the LLM. Similar steps are often taken in other data extraction approaches, and machine learning in general [35, 41, 58]. Specifically, as the highest scoring sentences are being manually verified, a new training sets consisting of true positive and true negative examples is built. Since the precision of results of Step 1 is typically around 50% at 90% recall (see Fig. 3 (d)), the created sets are typically close to equal in size. The human labeled sets consist of positive cases, which represent true positives from Step 1, and negative cases, consisting of false positives from Step 1. The latter are the most valuable counter-examples for the negative training set, as these are the sentences the easiest for the model to confuse for positives. If after reaching the desired amount of verified positive sentences the corresponding set of negative sentences is smaller, it may be complemented with random sentences from the analyzed papers (the exceeding majority of which are negative). Fig. 4 shows how the classification model improves when fine-tuned on datasets of increasing size. A detailed analysis of that figure is present in Sec. III, where a conclusion is made that after around 100 positive sentences for the quicker learning models such as GPT-3 davinci/GPT-3.5 or bart, we start to observe diminishing returns with this human labeled dataset size increase and it may not be worth spending more human time on obtaining more examples. Therefore, we recommend to perform the manual verification of the zero-shot classification until 100 positive sentences (and a corresponding 100 negative - made easy due to the  50% precision) are obtained, a number easy to remember and satisfactory for an efficient fine-tuning dataset.
This steps usually takes no more than 30 minutes for approximately 100 sentences - each sentence has to be classified only in a binary fashion which is a very simple task and takes just a few seconds per sentence. The classification is as simple and straightforward as assigning $1$ for positive and 00 to each sentence in a spreadsheet. The fine-tuning itself, for the small locally hosted models (bart and DeBERTaV3), takes around 30 minutes on an older workstation CPU (Intel(R) Xeon(R) CPU E5-2670), 20 minutes on a modern laptop CPU (Intel(R) Core(TM) i9-9880H), and can be reduced to just a few minutes if GPUs are used. The OpenAI models are fine-tuned on external OpenAI servers in less than 30 minutes and do not require any local resources. After this step is performed and the sentences are once again reclassified using the now fine-tuned model the precision and recall are greatly improved, as can be seen in Fig. 4.
@@
Data structurization (template filling). In this step, extraction of the structured data is performed. Here, by structured data we mean the full information necessary to provide a datapoint: the name of the material/system, the numerical value of the given property, its unit, and in some cases an additional detail, such as the temperature at which the value was obtained. At the same time as the data is extracted the precision of the result is improved to perfect (or near perfect, depending on the expertise and accuracy of the human supervising this step), by simply ignoring the false positive sentences left over from previous steps. The result of this step is a final, curated structured database.
@@
The user will typically perform this step by first ranking the sentences by their probability of being relevant (classification scores in the case of small LMs, bart and DeBERTaV3, or log probabilites in the case of GPT-3), which is the output from Step 1 (or Step 2 if performed), and start reviewing the list at the top, working down until they decide to stop. As the user works through the results in that fashion, they traverse down the precision recall curve (PRC) (see. Figs. 3 (a) and 4 (a)). While the recall is impossible to assess without knowing the ground truth, the user is fully aware of the precision of the data they have already analyzed, therefore using the PRC they can estimate the recall and decide to stop when a desired recall is reached (with the assumption that the PRC are similar to those shown in Figs. 3 (a) and 4 (a)). For best models, reaching recall of around 90% (close to that of a fully manual data curation) without performing the optional Step 2 happens for a precision close to 60%, while for a fine-tuned model (with Step 2), for a precision over 80%. It is entirely up to the user to decide the quality they require from their database, and the quality of the results will be proportional to the amount of time spent in this step. Recall of 90% seems to be a reasonable value to stop the process, as the precision sharply drops for higher values, which diminishes returns for the human time involved. However, this behavior may vary depending on the case, which will be discussed further in the Sec. III.
@@
In general, human assisted data structurization, even when only the sentences containing the relevant data are given, may be a tedious and time consuming task. However, at this point it is the only method that can guarantee an almost perfect precision. For an inexperienced user, extracting one datapoint from a given sentence and its surrounding context fully by hand may take as long as 30 seconds, depending on the complexity of the property being analyzed and how it is typically expressed in research papers. Considering this, only relatively modest sized databases are reasonable to create. However, with experience, this time quickly reduces as the user gets used to the process. In addition, more experienced users may employ simple computer codes, e.g. based on regular expressions, which would preselect possible candidates for values and units, reducing the time significantly. In the longer term, it is likely the NLP tools will help automate this data structurization step. Some models, like GPT4, offer structured format output, such as json, which may be used to assist the final data extraction step. However, they do not do this very effectively at present without either human supervision or a major effort to tune them. For example, GPT-3/3.5/4 is capable of parsing unstructured data in a zero-shot fashion, with no need for retraining. In the case of our bulk modulus sentences dataset we found that in over 60% of cases GPT-3/3.5/4 is capable of correctly providing the entire data entry for a given property (name of the material/system, value, unit), and an incomplete datapoint (wrong material/system, but correct value and unit) in over 95% of cases. The only drawback that prohibits a full automation of this step with a LLM is the inability to automatically and unambiguously distinguish between correct and incorrect extracted datapoints. Even though the model does not tend to make up (hallucinate) data, it sometimes provides an incomplete or inaccurate extraction (e.g. ”alloy” instead of ”AlCu alloy” for the material, or ”100” instead of ”greater than />/>/ > 100” for the value, etc.). However, human assistance in determining whether the data has been structurized properly, and in case it was not, fixing it by hand, can easily remedy that problem. Since almost all values and units are extracted correctly, and only less than half of the material names require fixing, using a LLM approach greatly reduces the human time and effort required for data structurization. Using an LLM we found that the average human time required to extract each good datapoint was reduced to under 10 seconds, keeping the same, almost perfect precision. Thus, and NLP-assisted data structurization, while still a tedious process, enables one create databases of up to around 1000 entries (more or less, depending on the users predisposition to and efficiency at repetitive tasks), in one workday. This timing includes the whole process, beginning (Step 0) to end (structurized database after Step 3), although almost all the human time is spent in Step 3.